OpenStudy (anonymous):

solve for x: 1/2 + 4/2x = x+4/10

OpenStudy (anonymous):

Is t (x+4)/10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is t (4/2)x or 4/(2x)?

OpenStudy (anonymous):

1/2 + 4/2x = x+4/10=(x+4)/10

OpenStudy (anonymous):